# Preference-Guided Diffusion for Multi-Objective Offline Optimization

### v1

@@ Abstract
footnotetext: Author email: {yashas.annadani, stefan.bauer}@helmholtz-munich.de, {syrineb, ermon, bengelhardt}@stanford.edu
@@ 1 Introduction
Several design problems in science and engineering require optimizing a black-box, expensive-to-evaluate function. For example, in antibiotic drug discovery, the goal is to identify novel molecules with high antibacterial activity [31]. This can be formulated as a single-objective optimization problem. However, in practice, most real-world design challenges involve balancing multiple conflicting objectives. For instance, in drug discovery, besides maximizing antibacterial activity, we also aim to minimize toxicity and production costs [30]. This constitutes a multi-objective experimental design problem.
@@
Prior work in both single and multi-objective optimization (MOO) has largely focused on adaptive experimental design using online methods such as Bayesian optimization [25]. These approaches rely on training surrogate models for each objective function and designing acquisition functions that are typically optimized via gradient-based techniques [3] or evolutionary algorithms to determine the next candidate for evaluation. This process is iteratively repeated to optimize the objectives. However, in many real-world applications, sequential evaluations—where inputs are tested one at a time or in small batches—are impractical. In some cases, we have only a single opportunity to evaluate the function, and, we must allocate the entire evaluation budget efficiently.
@@
For example, in drug design, scientists cannot test molecules one by one in wet lab experiments due to the high cost, slow turnaround, and the inherently parallelizable nature of the process [30]. Instead, it is common to evaluate all candidate molecules in a single batch. This setting is referred to as offline black-box optimization [33]. While recent work has explored offline optimization in the single-objective setting [38, 18], where extensive prior data is leveraged to model the objective function and identify potential optima, the MO case remains relatively underexplored.
@@
Offline black-box optimization presents distinct challenges compared to traditional online optimization. Since the algorithm cannot iteratively refine the learned model using newly acquired data, it must effectively leverage the available dataset to generalize beyond observed data points. This is particularly challenging because the true optima are often expected to lie outside the existing dataset, requiring robust extrapolation. Additionally, the goal is not merely to identify data points with high function values but to find solutions that satisfy a well-defined notion of optimality, such as Pareto optimality, in multi-objective settings.
@@
Prior work in single-objective offline optimization can generally be categorized into two main approaches. Forward approaches attempt to mimic strategies used in online optimization while leveraging offline data [33]. These methods train a surrogate model of the objective function and optimize it using gradient-based techniques to propose a set of promising inputs for evaluation. While effective when the search space is well-defined, forward methods rely on having a known set of candidate inputs to evaluate. In contrast, inverse approaches use generative models to learn an inverse mapping from function values to inputs, enabling the generation of new candidates with potentially high objective values [20, 5, 13, 19]. This distinction is critical in many real-world science and engineering problems where the optimal inputs are not known in advance. For example, in chemistry, if the goal is to evaluate a known molecule, surrogate models are effective in predicting its properties. However, if the goal is to discover entirely new molecules with desired properties, inverse methods are essential, as they directly generate novel candidates rather than selecting from a predefined space. Some generative modeling approaches also draw inspiration from online methods by constructing synthetic optimization trajectories from offline data, aiming to generate new optimal points by extrapolating from the learned trajectories [18].
Multi-objective offline optimization introduces additional challenges beyond those encountered in the single-objective setting. In single-objective optimization, the goal is simply to maximize (or minimize) a function value. However, in the multi-objective case, optimality is defined in terms of Pareto optimality, which seeks the best trade-offs among competing objectives. A solution is considered Pareto optimal in the multi-objective setting when no other solution in the search space dominates the Pareto optimal solution across all objectives, meaning that improving one objective would necessarily degrade another [4]. The Pareto front includes the objective values for all Pareto-optimal points (the Pareto set).
@@
Beyond identifying Pareto-optimal solutions, another critical challenge is ensuring diversity on the Pareto front. A well-structured Pareto front should provide solutions spread across different regions of the objective space, representing a broad range of Pareto-optimal designs. If all high-performing solutions are concentrated in a narrow region of the Pareto front, the optimization process fails to capture the full spectrum of desirable trade-offs, limiting its usefulness in real-world decision-making. This issue of diversity is well recognized even in online multi-objective optimization [1], where iterative refinement can be used to adjust the design choices. In the offline setting, however, ensuring both optimality and diversity becomes even more challenging, as the algorithm must infer these solutions solely from pre-existing data without the ability to iteratively refine its choices.
@@
Xue et al. [37] has recently explored benchmarking offline multi-objective optimization (MOO) by building offline datasets for a variety of MOO benchmarks and proposing several potential algorithms that can provide simple solutions. Their work extends some offline single-objective optimization (SOO) approaches to the MOO setting, such as fitting surrogate models to the offline data and optimizing over the surrogates using evolutionary algorithms. However, this benchmarking paper did not tackle or suggest any approach on how to use generative models for offline MOO.
In this work, we propose a novel algorithm that leverages diffusion models for offline MOO. Our approach formulates the problem as an inverse problem, using the generative capabilities of diffusion models to produce diverse and high-quality candidate solutions that extend beyond the regions of space covered by the training data. Diffusion models have demonstrated a strong ability to generate novel samples distinct from the training data [11], making them well-suited for this task.
To incorporate Pareto optimality into the generation process, we introduce a preference-based classifier that guides the diffusion model towards Pareto-optimal solutions. This classifier is trained to compare two candidate designs and determine which one is more likely to dominate the other in the objective space. By integrating this preference-based guidance into the sampling process, we steer the generative model toward regions containing Pareto-optimal solutions.
Additionally, we address the diversity challenge in the Pareto front generation. Rather than merely identifying high-function-value designs, we train the classifier to favor solutions that are not only optimal but also well-distributed across the objective space. To achieve this, we incorporate a crowding distance-based criterion into the preference pairs used for training. Solutions with higher crowding distance, indicating greater separation from neighboring points, are considered more diverse and are prioritized in the optimization process. This ensures that the resulting Pareto front is not overly concentrated in a limited region but instead captures a wide range of trade-offs between objectives.
@@ 2.1 Offline Multi-Objective Optimization
where $f_{i}:{\mathcal{X}}^{d}\mapsto{\mathbb{R}}$ is an unknown and expensive to evaluate objective. For most practical problems, the objectives are not simultaneously optimizable by a single design. Hence, the goal instead is to find the set of designs that are Pareto optimal.
@@ 2.2 Diffusion Models
Diffusion models [26, 14, 27] are a class of generative models defined by a Markov chain that sequentially adds noise to data samples and then learning to denoise it by reversing the Markov chain. In this work, we follow the Denoising Diffusion Probabilistic Models (DDPM) [14] approach, which we summarize here. Given a sample ${\bm{x}}_{0}\sim q({\bm{x}})$, a time-dependent forward noising process is defined as:
@@
The reverse process variance is set to be the same as the forward process variance at time $t$, i.e., $\Sigma_{\theta}({\bm{x}}_{t})\coloneqq\beta_{t}{\mathbb{I}}_{d}$. The denoising model can be trained with mean squared error loss:
@@ 2.3 Classifier Guidance
If label ${\bm{y}}$ corresponding to each sample ${\bm{x}}_{0}$ is available, then classifier guidance allows one to generate new samples from a trained diffusion model specific to a desired label. To do this, classifier guidance [11] trains an additional time-dependent classifier of the input $p_{\phi}({\bm{y}}\mid{\bm{x}}_{t},t)$. Along with the trained denoising model $\epsilon_{\theta}({\bm{x}}_{t},t)$, a new conditional sample can be generated by first sampling $\tilde{\bm{x}}_{T}\sim{\mathcal{N}}(0,{\mathbb{I}}_{d})$ and then, from $t=T$ to $t=1$, autoregressively sampling:
$\tilde{\bm{x}}_{t-1}\mid{\bm{y}},\tilde{\bm{x}}_{t}\sim{\mathcal{N}}({\bm{x}}_{t-1};\mu_{\theta}(\tilde{\bm{x}}_{t})+w\beta_{t}\nabla_{\tilde{\bm{x}}_{t}}\log p_{\phi}({\bm{y}}\mid\tilde{\bm{x}}_{t},t),\beta_{t}{\mathbb{I}}_{d}),$ (6)
where $w$ is the guidance strength. Classifier guidance has been successfully used in various domains including computer vision and audio to generate samples from a certain class or label. However, their utility is less explored in black-box optimization. In this work, we use classifier guidance to generate samples that approximate the optimal Pareto sets.
@@ 3 Related Work
Online Multi-Objective Black-Box Optimization
Existing work in adaptive experimental design has addressed the multi-objective optimization problem primarily in an online fashion, where solutions are refined iteratively based on newly acquired data [1, 4]. While research on multi-objective optimization is less extensive than in the single-objective setting, several approaches have successfully tackled the problem sequentially.
One of the most prominent methods is Bayesian optimization (BO), which typically consists of three main components: a surrogate model, an acquisition function, and an optimizer. The surrogate model approximates the objective functions based on previously observed data, while the acquisition function, built using the surrogate model, is optimized to determine the next input to evaluate [25]. Most BO methods rely on Gaussian processes [36], particularly in data-scarce scenarios where no prior data is available. Some approaches simplify the multi-objective problem by reducing it to a single-objective formulation using scalarization techniques, such as linear scalarization or Chebyshev scalarization [16]. More advanced methods have leveraged expected hypervolume improvement (EHVI) [12] or information gain-based acquisition functions [4] to better navigate the trade-offs between competing objectives.
Recent advances in BO have also introduced batch selection strategies improving efficiency in parallelizable experiments [9]. However, while most of these approaches focus on identifying Pareto-optimal solutions, a few methods including Konakovic Lukovic et al. [17], Ahmadianshalchi et al. [1] explicitly address the diversity of the Pareto front. Instead, the majority of existing techniques prioritize Pareto dominance while neglecting the distribution of solutions across the objective space.
Beyond small models, some approaches have explored the use of neural networks to extend online MOO techniques by leveraging existing data. These methods often incorporate generative models, such as variational autoencoders (VAEs), combined with Gaussian processes over the latent space of the VAE [28]. However, these techniques inherit VAE’s well-known challenges, including posterior collapse and unidentifiability, which can limit their effectiveness in practical optimization settings.
Offline Single Objective Black-Box Optimization
Several approaches have been proposed to tackle single-objective offline optimization. Forward approaches, such as Trabucco et al. [33], Yu et al. [38], train a surrogate model to approximate the objective function and then optimize this model to identify a set of inputs with potentially optimal performance. Other methods [20, 5, 13] adopt generative adversarial networks (GANs) with conditional generation to learn an inverse mapping from the function space to the input space. However, these approaches inherit the well-known challenges of GANs, including unstable training dynamics and mode collapse, which can hinder their reliability in practical applications.
To address the limitations of standard forward and inverse approaches, Chemingui et al. [6], Krishnamoorthy et al. [18] proposed methods that mimic online optimization by constructing synthetic optimization trajectories from offline data. This approach trains a sequence-based model to generate new candidate solutions by extrapolating from learned optimization paths.
More recently, Krishnamoorthy et al. [19] introduced a diffusion-based optimization framework, modifying the diffusion model’s loss function to incorporate weighted importance terms. This adjustment encourages the model to prioritize inputs with higher function values while penalizing those with lower values, effectively guiding the generative process toward promising regions of the input space.
Additionally, Trabucco et al. [34] proposed a benchmarking framework to standardize and facilitate the evaluation of offline optimization algorithms, providing a consistent platform for comparing different approaches.
Offline Multi-Objective Black-Box Optimization
There is very limited prior work on offline multi-objective optimization (MOO). Xue et al. [37] recently explored this area by introducing a benchmarking framework, where they constructed offline datasets for various MOO benchmark problems and proposed several baseline algorithms. Their approach primarily involved extending forward techniques from offline single-objective optimization (SOO) to the multi-objective setting, such as fitting surrogate models to offline data and optimizing over these surrogates using evolutionary algorithms. However, this work did not explore or propose methods for leveraging generative models for inverse problems in offline MOO.
A concurrent study by Yuan et al. [40] introduces a flow-based generative model for offline MOO. Their approach scalarizes the objective functions by incorporating a weighting scheme directly into the loss function of the generative model. While this method provides an inverse method perspective, it does not address the diversity challenge in Pareto front generation, which is a critical aspect of multi-objective optimization.
4 Method: Preference-Guided Diffusion for Offline Multi-Objective Optimization
@@
We present a new effective approach for offline MOO by using classifier guidance to generate samples from Pareto optimal sets with a diffusion model trained on offline data. Our approach does not require training individual surrogate models for each objective. It relies on an inverse strategy while ensuring the ability to generate diverse samples from the Pareto optimal set. We refer to our method as Preference-Guided Diffusion for Multi-objective Offline optimization (PGD-MOO)
@@
Let ${\bm{x}}\in{\mathcal{X}}^{d}\subseteq{\mathbb{R}}^{d}$ be any $d$-dimensional design with corresponding objective values $y_{i}=f_{i}({\bm{x}})$ defined by unknown and expensive to evaluate functions $f_{i}:{\mathcal{X}}^{d}\mapsto{\mathbb{R}}$. Let ${\bm{y}}=[y_{1},\dots,y_{m}]^{T}$ be the vector of objective values for an $m$-objective problem. In offline MOO, we have access to a dataset ${\mathcal{D}}:=\{{\bm{x}}^{(i)},{\bm{y}}^{(i)}\}_{i=1}^{N}$ of $N$ previously evaluated design-objective pairs. Given ${\mathcal{D}}$, the goal is to generate designs ${\bm{x}}^{*}$ from the unknown optimal Pareto set.
@@
While diffusion models capture the distribution over data $p({\bm{x}})$, in offline MOO, we are often interested in samples that lie outside the training data, closer to the Pareto front. This motivates the use of classifier guidance. Directly using classifier guidance in diffusion models usually involves training surrogate models for each objective, which often requires scalarization and hence can be suboptimal. We propose to use preference-based guidance to capture Pareto dominance relations between data points.
@@
Figure 1: Generalization of the preference model on regions unseen in the training data on the ZDT2 task [45]. The preference model gives good prediction of Pareto dominance between reference design (in black) with other designs. The figure on the right is a zoomed-in version of the left, excluding the training data (in blue).

### v2

@@ 1 Introduction
Several design problems in science and engineering require optimizing a black-box, expensive-to-evaluate function. For example, in antibiotic drug discovery, the goal is to identify novel molecules with high antibacterial activity [swanson2024generative]. This can be formulated as a single-objective optimization problem. However, in practice, most real-world design challenges involve balancing multiple conflicting objectives. For example, in drug discovery, in addition to maximizing antibacterial activity, we also want to minimize toxicity and production costs [stokes2020deep]. This constitutes a multi-objective experimental design problem.
@@
Prior work in both single and multi-objective optimization (MOO) has largely focused on adaptive experimental design using online methods such as Bayesian optimization [shahriari2015taking]. These approaches rely on training surrogate models for each objective function and designing acquisition functions that are typically optimized via gradient-based techniques [balandat2020botorch] or evolutionary algorithms to determine the next candidate for evaluation. This process is iteratively repeated to optimize the objectives. However, in many real-world applications, sequential evaluations—where inputs are tested one at a time or in small batches—are impractical. In some cases, we have only a single opportunity to evaluate the function, and we must allocate the entire evaluation budget efficiently.
@@
For example, in drug design, scientists cannot test molecules one by one in wet lab experiments due to the high cost, slow turnaround, and the inherently parallelizable nature of the process [stokes2020deep]. Instead, it is common to evaluate all candidate molecules in a single batch. This setting is referred to as offline black-box optimization [trabucco2021conservative]. While recent work has explored offline optimization in the single-objective setting [yu2021roma, krishnamoorthy2022generative], where extensive prior data is leveraged to model the objective function and identify potential optima, the MO case remains relatively underexplored.
@@
Offline black-box optimization presents distinct challenges compared to traditional online optimization. Since the algorithm cannot iteratively refine the learned model using newly acquired data, it must effectively leverage the available dataset to generalize beyond observed data points. This is particularly challenging because the true optima are often expected to lie outside the existing dataset, requiring robust extrapolation. Additionally, the goal is not merely to identify data points with high function values but to find solutions that satisfy a well-defined notion of optimality, such as the Pareto optimality.
@@
Prior work in single-objective offline optimization can generally be categorized into two main approaches: forward and inverse methods. Forward approaches attempt to mimic strategies used in online optimization while leveraging offline data [trabucco2021conservative]. These methods train a surrogate model of the objective function and optimize it using gradient-based techniques to propose a set of promising inputs for evaluation. They are effective when the search space is well-defined. In contrast, inverse approaches use generative models to learn an inverse mapping from function values to inputs, enabling the generation of new candidates with potentially high objective values [kumar2020model, brookes2019conditioning, fannjiang2020autofocused, krishnamoorthy2023diffusion]. This distinction is critical when optimal inputs are not known in advance. For example, in chemistry, if the goal is to evaluate a known molecule, surrogate models are effective in predicting its properties. However, if the goal is to discover entirely new molecules with desired properties, inverse methods are essential, as they directly generate novel candidates rather than selecting from a predefined space. Some generative modeling approaches also draw inspiration from online methods by constructing synthetic optimization trajectories from offline data, aiming to generate new optimal points by extrapolating from the learned trajectories [krishnamoorthy2022generative].
@@
Multi-objective offline optimization introduces additional challenges beyond those encountered in the single-objective setting. In single-objective optimization, the goal is simply to maximize (or minimize) a function value. However, in the multi-objective case, we seek to achieve best trade-offs among competing objectives, which is formalized by Pareto optimality. Beyond identifying Pareto-optimal solutions, another critical challenge is ensuring diversity on the Pareto front. A well-structured Pareto front should provide solutions spread across different regions of the objective space, representing a broad range of Pareto-optimal designs. Although the problem of ensuring diversity of solutions is addressed in online MOO [ahmadianshalchi2024pareto], ensuring both optimality and diversity is challenging in the offline setting as the algorithm must infer these solutions solely from existing data.
@@
xue2024offline has recently explored benchmarking offline multi-objective optimization (MOO) by building offline datasets for a variety of MOO benchmarks and proposing several potential algorithms. Their work extends some offline single-objective optimization (SOO) approaches to the MOO setting, such as fitting surrogate models to the offline data and optimizing over the surrogates using evolutionary algorithms. However, this work has not explored the possibility of using generative models for offline MOO.
Our contributions are:
  • We formulate offline MOO as an inverse problem, and propose a novel algorithm based on diffusion models with preference-based classifier guidance. This classifier is trained to compare two candidate designs and determine which is more likely to dominate the other in the objective space.
  • We address the diversity challenge in the offline MOO setting. Rather than just identifying high-value designs, we train the preference model to favor solutions that are not only optimal but also well-distributed across the objective space.
  • We demonstrate through extensive experiments on both synthetic and real-world settings that our approach performs better than other inverse methods in terms of convergence to the Pareto front and the diversity of obtained solutions.
  • Among inverse techniques, our approach consistently achieves the strongest results; making it a competitive choice in cases where forward methods are not feasible. Even when forward methods are applicable, our algorithm remains competitive while maintaining the benefits of generative exploration.
Figure 1: A schematic representation of the proposed preference guided diffusion approach for offline multi-objective optimization.
@@ 2.1 Offline Multi-Objective Optimization
where $f_{i}:{\mathcal{X}}^{d}\mapsto{\mathbb{R}}$ is an unknown and expensive-to-evaluate objective. For most practical problems, the objectives are not simultaneously optimizable by a single design. Hence, the goal instead is to find the set of designs that are Pareto optimal.
@@ 2.2 Diffusion Models
Diffusion models [sohl2015deep, ho2020denoising, song2020score] are a class of generative models defined by a Markov chain that sequentially adds noise to data samples and then learns to denoise from white noise (i.e., generate a sample) by reversing the Markov chain. In this work, we follow the Denoising Diffusion Probabilistic Models (DDPM) [ho2020denoising] approach, which we summarize here. Given a sample ${\bm{x}}_{0}\sim q({\bm{x}})$, a time-dependent forward noising process is defined as:
@@
The reverse process variance is set to be the same as the forward process variance at time $t$, i.e., $\Sigma_{\theta}({\bm{x}}_{t})\coloneqq\beta_{t}{\mathbb{I}}_{d}$. The denoising model can be trained with mean squared error (MSE) loss:
@@ 2.3 Classifier Guidance
If label $\bm{o}$} corresponding to each sample ${\bm{x}}_{0}$ is available, then classifier guidance allows one to generate new samples from a trained diffusion model specific to a desired label. To do this, classifier guidance [dhariwal2021diffusion] trains an additional time-dependent classifier of the input $p_{\phi}({\bm{o}}\mid{\bm{x}}_{t},t)$. Along with the trained denoising model $\epsilon_{\theta}({\bm{x}}_{t},t)$, a new conditional sample can be generated by first sampling $\tilde{\bm{x}}_{T}\sim{\mathcal{N}}(0,{\mathbb{I}}_{d})$ and then, from $t=T$ to $t=1$, autoregressively sampling:
$\tilde{\bm{x}}_{t-1}\mid{\bm{o}},\tilde{\bm{x}}_{t}\sim{\mathcal{N}}({\bm{x}}_{t-1};\mu_{\theta}(\tilde{\bm{x}}_{t})+w\beta_{t}\nabla_{\tilde{\bm{x}}_{t}}\log p_{\phi}({\bm{o}}\mid\tilde{\bm{x}}_{t},t),\beta_{t}{\mathbb{I}}_{d}),$ (6)
where $w$ is the guidance strength. In this work, we use classifier guidance to generate samples that approximate the optimal Pareto sets where the classifier is a preference model that predicts the dominance of one input over the other.
@@ 3 Related Work
Online Multi-Objective Black-Box Optimization:
Adaptive experimental design has primarily explored multi-objective optimization (MOO) in an online or sequential fashion, where solutions are iteratively refined as new data arrives [ahmadianshalchi2024pareto, belakaria2021output]. Although less studied than single-objective optimization, several sequential approaches have proven effective. The most established is Bayesian optimization (BO) [jones1998efficient, shahriari2015taking], which typically uses Gaussian processes [williams2006gaussian] to model objectives, especially in data-scarce regimes. Multi-objective extensions of BO often reduce the problem via scalarization [knowles2006parego] or employ acquisition functions such as expected hypervolume improvement (EHVI) [emmerich2008computation] and information gain [belakaria2021output] to balance trade-offs across objectives. Recent work also introduced batch selection strategies [daulton2020differentiable], though most methods emphasize Pareto dominance over diversity. A few exceptions, such as konakovic2020diversity and ahmadianshalchi2024pareto, explicitly promote Pareto front diversity. Beyond traditional BO, neural approaches extend online MOO using generative models like variational autoencoders (VAEs) combined with Gaussian processes over the latent space [stanton2022accelerating]. However, these methods inherit VAE limitations—posterior collapse and non-identifiability—that restrict their robustness in practical optimization tasks.
Offline Single-Objective Black-Box Optimization:
Offline single-objective optimization focuses on leveraging existing datasets without additional data collection. Forward methods [trabucco2021conservative, yu2021roma] train surrogate models to approximate the objective function and then optimize these surrogates for high-performing inputs. Inverse methods instead employ conditional GANs [kumar2020model, brookes2019conditioning, fannjiang2020autofocused] to learn mappings from function outputs back to inputs. To bridge forward and inverse strategies, chemingui2024offline, krishnamoorthy2022generative proposed hybrid methods that synthesize pseudo-optimization trajectories from offline data, mimicking online behavior. Diffusion-based frameworks [krishnamoorthy2023diffusion] adjust diffusion losses with weighted importance terms, while classifier-guided models [chen2024robust] bias generation toward globally optimal designs. kim2025offline provides a comprehensive review of model-based offline optimization methods, summarizing progress across forward, inverse, and hybrid approaches. Additionally, trabucco2022design established a standardized benchmarking suite for consistent evaluation of offline optimization algorithms.
Offline Multi-Objective Black-Box Optimization:
Research on offline multi-objective optimization (MOO) remains limited. xue2024offline recently proposed a benchmarking framework with offline datasets and baseline algorithms across MOO benchmarks. Their work mainly extends forward-style single-objective methods by fitting surrogate models to offline data and optimizing them via evolutionary algorithms but does not explore generative or inverse techniques. Concurrently, yuan2024paretoflow introduced ParetoFlow, a flow-based generative model for offline MOO that embeds objective weighting directly into the loss function, effectively scalarizing multiple objectives. While this inverse approach is promising, it does not explicitly address Pareto front diversity—an essential factor for capturing representative trade-offs in MOO.
4 Preference-Guided Diffusion for Offline Multi-Objective Optimization
@@
We present a new effective approach for offline MOO by using classifier guidance to generate samples from Pareto optimal sets with a diffusion model trained on offline data. Our approach does not require training individual surrogate models for each objective. It relies on an inverse strategy while ensuring the ability to generate diverse samples from the Pareto optimal set. We refer to our method as preference-guided diffusion for multi-objective offline optimization (PGD-MOO).
@@
Let ${\bm{x}}\in{\mathcal{X}}^{d}\subseteq{\mathbb{R}}^{d}$ be any $d$-dimensional design with corresponding objective values $y_{i}=f_{i}({\bm{x}})$ defined by unknown and expensive-to-evaluate functions $f_{i}:{\mathcal{X}}^{d}\mapsto{\mathbb{R}}$. Let ${\bm{y}}=[y_{1},\dots,y_{m}]^{T}$ be the vector of objective values for an $m$-objective problem. In offline MOO, we have access to a dataset ${\mathcal{D}}:=\{{\bm{x}}^{(i)},{\bm{y}}^{(i)}\}_{i=1}^{N}$ of $N$ previously evaluated design-objective pairs. Given $\mathcal{D}$}, the goal is to generate designs $\bm{x}$$^{*}$ from the unknown optimal Pareto set.
@@
While diffusion models capture the distribution over data $p({\bm{x}})$, in offline MOO, we are often interested in samples that lie outside the training data, closer to the Pareto front. This motivates the use of classifier guidance. Directly using classifier guidance in diffusion models involves training surrogate models for each objective, which often requires scalarization and hence can be suboptimal. We propose to use preference-based guidance to capture Pareto dominance relations between data points.
@@
Figure 2: Generalization of the preference model on regions unseen in the training data on the ZDT2 task [zitzler2000comparison]. The preference model gives good prediction of Pareto dominance between the reference design (in black) with other designs (in pink and green). Pink indicates that the preference model predicts these designs to be dominated by the reference design and green indicates that these designs are predicted as dominating the reference design. The figure on the right is a zoomed-in version of the left, excluding the training data (in blue).